OpenStudy (anonymous):

s e n d + m o r e ----------------- m o n e y Each letter represent an integer between 0 and 9 Two different letters represent two different numbers. m o n e y = s e n d + m o r e find all the value of all the letters. There is a story about this problem. A student wanted money for her father who was good mathematician. She sends her father a letter containing s e n d + m o r e ----------------- m o n e y The father solved the problems and sent her $ m o n e y.

OpenStudy (experimentx):

m=1

OpenStudy (anonymous):

Great start

OpenStudy (experimentx):

lol

OpenStudy (anonymous):

hello, again doc... great puzzle btw... :)

OpenStudy (across):

if m=1 then s>=8

OpenStudy (experimentx):

s can be 8 or 9 ... lol i was thinking the same

OpenStudy (experimentx):

o must be 0

OpenStudy (zepp):

If o = 0 then E+1 = N

OpenStudy (experimentx):

s must be 9

OpenStudy (experimentx):

n+r = 10 + e e + 1 = n => r = 9 <--- something wrong gone here!!

OpenStudy (experimentx):

n+r +1= 10 + e e + 1 = n => r = 8

OpenStudy (experimentx):

n> 2

OpenStudy (experimentx):

gone 9,8,0,1

OpenStudy (zepp):

If we replace N by E+1 in the NRE row, E+1+R = E

OpenStudy (zepp):

That makes me to believe that D+E > 10

OpenStudy (zepp):

and N+R > 10 too

OpenStudy (zepp):

Did a test there

OpenStudy (across):

I wrote a short program and found that there are 25 different answers to this problem. :/

OpenStudy (experimentx):

i'm out of clues!!

OpenStudy (zepp):